### v1

# Fast nonparametric inference of network backbones for graph sparsification

Abstract
A network backbone provides a useful sparse representation of a weighted network by keeping only its most important links, permitting a range of computational speedups and simplifying complex network visualizations. There are many possible criteria for a link to be considered important, and hence many methods have been developed for the task of network backboning for graph sparsification. These methods can be classified as global or local in nature depending on whether they evaluate the importance of an edge in the context of the whole network or an individual node neighborhood. A key limitation of existing network backboning methods is that they either artificially restrict the topology of the backbone to take a specific form (e.g. a tree) or they require the specification of a free parameter (e.g. a significance level) that determines the number of edges to keep in the backbone. Here we develop a completely nonparametric framework for inferring the backbone of a weighted network that overcomes these limitations by automatically selecting the optimal number of edges to retain in the backbone using the Minimum Description Length (MDL) principle from information theory. We develop two encoding schemes that serve as objective functions for global and local network backbones, as well as efficient optimization algorithms to identify the optimal backbones according to these objectives with runtime complexity log-linear in the number of edges. We show that the proposed framework is generalizable to any discrete weight distribution on the edges using a maximum a posteriori (MAP) estimation procedure with an asymptotically equivalent Bayesian generative model of the backbone. We compare the proposed method with existing methods in a range of tasks on real and synthetic networks.
I Introduction
In a range of practical applications including the simulation of epidemics or information cascades mercier2022effective ; mathioudakis2011sparsification and the visualization of large networks imre2020spectrum , it is often helpful to alleviate computational burden by reducing the number of edges in a network while preserving key properties of interest for computations. This task is known as graph sparsification spielman2008graph , and often involves identifying a network backbone consisting of a network subgraph in which important characteristics such as degree and strength distributions serrano2009extracting or shortest paths grady2012robust are preserved. As there are many possible objectives for sparsification, there have been many proposed methods for identifying backbones in networks neal2022backbone ; yassin2023evaluation . Some of these methods can be used to sparsify unweighted graphs, but they most often applied to weighted graphs in which edges with higher weight are more likely to be kept in the network backbone when the weight signals edge importance.
A number of existing methods for graph sparsification are stochastic, aiming to preserve properties such as network spectra with high probability when sampling edges based on their structural properties spielman2011spectral ; su2024generic ; john2017single . However, most network backboning methods are deterministic in nature and prune the network down to a single subgraph for analyses. Network backboning methods can generally be categorized as either global or local in nature yassin2023evaluation . Global backboning methods consider the whole structure of the network when considering whether or not to remove an edge, while local backboning methods only consider the neighborhood surrounding the nodes at one or both ends of the edge in question.
The simplest method for network backboning is a global thresholding procedure, in which all edges with weight less than some pre-defined threshold are dropped from the network. However, this has been shown to poorly preserve many structural properties of interest, especially surrounding nodes with generally weak ties serrano2009extracting . Another straightforward global backboning method is to compute the minimum (maximum) spanning tree with respect to weights transformed to represent some notion of distance (similarity) between nodes wu2006transport . Spanning trees have many useful applications but are naturally acyclic so cannot possibly preserve important network properties such as transitivity, clustering, or path lengths. The High Salience Skeleton (HSS) introduced in grady2012robust is a global backboning method that aims to identify edges critical for routing along shortest paths in the network. This method has the appealing properties of being principled and relatively insensitive to its only user-chosen threshold, but has a computational burden that is at least quadratic in the number of nodes in practice and often does not preserve the global connectivity of the graph (as we will see in Sec. III.2). Other global backboning methods tend to suffer from the same issue of scalability, as they often require the computation of shortest paths zhang2018extracting ; simas2021distance or community structure as input rajeh2022modularity .
Local backboning methods provide a scalable alternative to global methods by considering edges based only on nearby network structure, with the tradeoff that they may fail to preserve large-scale properties. One of the most popular methods for local network backboning is the Disparity Filter serrano2009extracting , which drops edges if their weight falls below a pre-specified significance level under a suitable null model relative to the other edges in a node’s neighborhood. The Polya Filter marcaccioli2019polya generalizes the disparity filter to a whole family of Polya urn null models, allowing for greater flexibility in modeling weight heterogeneity. A number of other statistical methods have also been considered that aim to remove edges that are deemed insignificant under a suitable null model tumminello2011statistically ; dianati2016unwinding ; casiraghi2017relational which may be distribution-free foti2011nonparametric . All of these methods have the requirement of a user-defined threshold or significance level that directly determines the number of edges present in the final backbone. While this approach can be advantageous for identifying multiple scales at which a network exhibits significant structural patterns serrano2009extracting , it also leaves arguably the most critical backboning choice up to the user and does not have a natural mechanism for selecting the single backbone representation to be used in an application.
Properly regularized Bayesian methods for network reconstruction have circumvented the need for fixing free parameters when inferring statistically significant network structure newman2018network ; peel2022statistical ; peixoto2024network . These methods are statistically principled and robust, but are usually not suitable for backboning applications because they often increase the number of edges in a network when there is strong evidence for edges being omitted during the measurement process. They also tend to be computationally expensive due to the need to search over edge configurations involving node pairs that do not have any observed edge, the number of which scales quadratically with the network size. There are fast approximations aimed to alleviate this issue and provide better asymptotic runtime scaling, peixoto2024scalable , although the resulting scaling is still worse than that of many backboning methods due to the inherently higher complexity of the reconstruction problem.
Here we develop a fast, completely parameter-free method for inferring the backbone of a weighted network based on the Minimum Description Length (MDL) principle from information theory rissanen1978modeling . Our description length objective computes the amount of information required to transmit the edge weights for a given network topology using a high weight backbone as an intermediate step. Minimizing over this encoding with an asymptotically optimal greedy procedure provides the backbone that best compresses the edge weight structure in the network, while automatically selecting for the number of edges to retain in the backbone. We develop two encodings using our framework—a global encoding and a local encoding—among which we can rigorously select an optimal backbone by comparing the final description length values. We show that the MDL framework is generalizable to any discrete edge weight distribution by reframing the problem as a Bayesian maximum a posteriori estimation procedure, allowing for a generative interpretation of the backbone formulation. Our method provides a flexible, principled, interpretable solution for network backboning that easily scales to large networks. We test our method on a range of real and synthetic network data, finding that it compares favorably against existing methods in a number of dimensions while eliminating the need for tuning any parameters.
II Methods
The method we derive here for network backboning is motivated by the Minimum Description Length (MDL) principle, which states that the optimal model for a dataset is the model that allows us to transmit the dataset using the fewest bits rissanen1978modeling . By considering both the information to transmit a dataset given a model and the information to transmit the model itself, the MDL principle provides a flexible framework for penalizing overly complex models to reduce the risk of overfitting. The MDL principle has been employed in the network context for a number of methods in the past, for the tasks of community detection peixoto2019bayesian ; kirkley2022spatial ; peixoto2023implicit , graph comparison coupette2021graph ; coupette2022differentially , graph summarization liu2018graph ; kirkley2023compressing , and for extracting structural regularities in dynamic networks and hypergraphs shah2015timecrunch ; young2021hypergraph ; kirkley2024identifying ; kirkley2024inference . In the backboning context, the MDL principle allows us to naturally promote sparsity in the inferred backbone while encouraging the backbone to carry the most significant weights in the network.
We consider two encodings for network weights that utilize a backbone to compress the weight information in the edges of the network. We first develop a global encoding and then apply similar reasoning to adapt this encoding to transmit node neighborhoods individually. We demonstrate that both encodings can be minimized using fast greedy procedures, allowing us to easily identify the MDL-optimal backbone configurations. We then demonstrate the equivalence of our MDL formulation with MAP estimation using a Bayesian generative model for the edge weights, allowing for generalization of our framework to a broader class of weight distributions on the edges.
II.1 Global Network Backboning Algorithm
Our problem considers an input network $G=\{(i,j,w_{ij})\}$, which we represent as a set of $E$ tuples $(i,j,w_{ij})$ that indicate an edge from node $i$ to node $j$ with weight $w_{ij}$. $G$ may have self-edges or multi-edges—multi-edges can be combined into a single edge weight—but the weights $w_{ij}$ must be positive and integer-valued. $G$ may be directed or undirected, and in the undirected case we simply consider both $(i,j,w_{ij})$ and $(j,i,w_{ij})$ to be in $G$, with the edge $(i,j,w_{ij})$ appearing in an undirected backbone if either $(i,j,w_{ij})$ or $(j,i,w_{ij})$ is retained in the directed backbone (as done in serrano2009extracting ). Bipartite networks can also be accommodated by our framework. We will let $W=\sum_{(i,j,w_{ij})\in G}w_{ij}$ denote the total edge weight in the weighted graph $G$. The goal of our network backboning method is to extract a subset $G^{(b)}\subseteq G$ of weighted edges that best compress the overall weight distribution in $G$. We call the subset $G^{(b)}$ the backbone of $G$, which has $E^{(b)}\leq E$ edges with a total weight of $W^{(b)}\leq W$.
To find the optimal backbone $G^{(b)}$ corresponding to an input network $G$, we consider transmitting the weight structure in $G$ to a receiver using $G^{(b)}$ as an intermediate step. We assume the edge topology of $G$ is known by the receiver and we must transmit the edge weight $w_{ij}$ corresponding to every edge $(i,j)$. This is the same assumption made implicitly by many network backboning methods serrano2009extracting ; marcaccioli2019polya ; dianati2016unwinding , which consider the statistical significance of edge weights while conditioning on the degrees of nodes.
The first transmission scheme we consider encodes the backbone $G^{(b)}$ and the remaining non-backbone edges $G\setminus G^{(b)}$ in a global manner, accounting for the distribution of weights across all edges when choosing to keep an edge in the backbone $G^{(b)}$. In this encoding we must first specify the total weight $W$ of the graph. Since the receiver knows the topology of $G$, they know the total number of edges $E$, and since the network must have positive integer-valued weights they also know that each edge has a weight of at least $1$. To specify the total weight $W$ we can thus specify the excess weight $W-E$ beyond the minimum weight of $E$ if all edges had weight $w_{ij}=1$. To specify $W-E$, we must send a bitstring of length approximately $\log(W-E+1)$ bits (the binary representation of the integer $W-E$), where $\log\equiv\log_{2}$. We then must specify the total weight $W^{(b)}$ of the backbone and the total number of edges $E^{(b)}$ in the backbone, which must fall in the ranges $[0,W]$ and $[0,E]$ respectively so require $\log(W+1)$ and $\log(E+1)$ bits respectively to specify.
The next step is to transmit which edges $(i,j)$ are in the backbone. There are $E$ possible edges to choose from, and the receiver now knows there are $E^{(b)}$ edges in the backbone, meaning there are (${E\choose E^{(b)}}$ possibilities for the backbone edges. Therefore to specify the particular configuration among these possibilities we need a bitstring of length approximately $\log{E\choose E^{(b)}}$. Since the weight of each edge $(i,j)$ in the backbone will satisfy $w_{ij}\geq 1$, the list of weights for these edges will take the form of a length-$E^{(b)}$ positive integer vector that sums to $W^{(b)}$. The number of such vectors is equal to (${W^{(b)}-1\choose E^{(b)}-1}$, which can be computed using the standard “stars and bars” argument from combinatorics feller1991introduction . Therefore specifying the backbone weights requires $\log{W^{(b)}-1\choose E^{(b)}-1}$ additional bits. Finally, we specify the non-backbone edges $G\setminus G^{(b)}$ using the same approach, which through an analogous argument requires $\log{W-W^{(b)}-1\choose E-E^{(b)}-1}$ bits.
Figure 1: Global and local MDL backbones for a small example network. (a) Schematic of the global backbone transmission objective of Eq. II.1, with the total weight and number of edges indicated for the backbone $G^{(b)}$ (left) and remaining edges (right). (b) Schematic of the local backbone transmission objective of Eq. 5, applied to out-neighborhoods $\partial_{i}$ for each node $i$. The total strength and degree are indicated for the backbone neighborhood $\partial^{(b)}_{i}$ around an example node $i$. Each method aims to transmit a backbone that parsimoniously captures the weight structure in the network $G$ according to an information encoding and learns an optimal number of backbone edges nonparametrically. The local method provides a more compressive backbone for this network, giving an inverse compression ratio (Eq. 6) of $\eta^{(\text{local})}=0.86$ versus $\eta^{(\text{global})}=0.88$ for the global backbone. The backboning methods are applicable to any network with integer-valued edge weights, including directed or undirected networks and unipartite or bipartite networks. The global and local backboning methods are also generalizable to a broader set of weight distributions on the edges using an equivalent Bayesian model, as discussed in Sec. II.3.
Putting this all together, we have that the description length of the network weights in $G$ using a backbone $G^{(b)}$ is given by
$\mathcal{L}^{(\text{global})}(G^{(b)})$ $=\log(W-E+1)+\log(W+1)$
$+\log(E+1)+\log{E\choose E^{(b)}}$ (1)
$+\log{W^{(b)}-1\choose E^{(b)}-1}+\log{W-W^{(b)}-1\choose E-E^{(b)}-1}.$
This encoding scheme is considered “global” because the description length couples all edges together under the total backbone weight $W^{(b)}$ and the backbone edge count $E^{(b)}$. By minimizing $\mathcal{L}^{(\text{global})}(G^{(b)})$ over all possible backbones $G^{(b)}$, we obtain the optimal backbone $G_{\text{global}}^{(b)}$ according to this global encoding scheme.
For a fixed number of edges $E^{(b)}$ in the backbone $G^{(b)}$, the description length in Eq. II.1 is monotonically decreasing in $W^{(b)}$ for $W,E\gg 1$. We can show this by computing
$\mathcal{L}^{(\text{global})}(W^{(b)}+1,E^{(b)})-\mathcal{L}^{(\text{global})}(W^{(b)},E^{(b)})$
$=\log{\frac{W_{b}}{W_{b}-E_{b}+1}\frac{(W-W_{b})-(E-E_{b})}{W-W_{b}-1}},$ (2)
which is less than or equal to zero for
$\frac{W^{(b)}}{W}\geq\frac{E^{(b)}}{E}+\frac{2W^{(b)}-W-E^{(b)}+1}{WE}.$ (3)
For $W,E\gg 1$, the second term on the RHS vanishes and we have that for $\frac{W^{(b)}}{W}\geq\frac{E^{(b)}}{E}$, the description length in Eq. II.1 is monotonically decreasing in $W^{(b)}$ for fixed $E^{(b)}$. Therefore, under this condition, we have that at any value of $E^{(b)}$ the best edge to add to the backbone $G^{(b)}$ is the remaining edge $(i,j)$ with the highest weight $w_{ij}$.
The above argument implies that a greedy edge addition procedure for constructing the backbone will be optimal for $W,E\gg 1$, since the condition $\frac{W^{(b)}}{W}\geq\frac{E^{(b)}}{E}$ must always hold under such a procedure—this condition is equivalent to the condition $\frac{W^{(b)}}{E^{(b)}}\geq\frac{W}{E}$, or that the average weight of the backbone is at least the average weight of the network. We can therefore approximately minimize the description length objective in Eq. II.1 in the regime $W,E\gg 1$ with the following simple algorithm:
  1. 1.
    Initialize an empty backbone $G^{(b)}=\{\}$ with $E^{(b)}=W^{(b)}=0$ and sort the edges $(i,j,w_{ij})\in G$ in decreasing order of weight $w_{ij}$. Initialize an array DL to store the description length value for each value of $E^{(b)}$. Compute $\text{DL}[E^{(b)}=0]$ using Eq. II.1.
  2. 2.
    Iterate through the ordered edges in $G$, and for each edge $(i,j,w_{ij})$: increment $E^{(b)}\mathrel{+}=1$ and $W^{(b)}\mathrel{+}=w_{ij}$ and compute $\text{DL}[E^{(b)}]$ using Eq. II.1.
  3. 3.
    After adding the last edge from $G$, identify the value $E_{\text{global}}^{(b)}=\text{argmin}[\text{DL}]$ for which the description length was minimized, and set the MDL-optimal backbone $G^{(b)}_{\text{global}}$ to be the first $E_{\text{global}}^{(b)}$ edges in the ordered $G$.
The bottleneck of this algorithm is simply the initial step of ordering the edges in $G$ by weight, so the overall runtime complexity is $O(E\log E)$—only marginally slower than constructing the edge set $G$ itself.
We can see that this optimization procedure will construct backbones $G^{(b)}$ that concentrate as much weight $W^{(b)}$ as possible on $E^{(b)}$ edges, up to some point at which it is no longer information theoretically preferable to continue adding edges to the backbone. The final backbone $G^{(b)}_{\text{global}}$ is the most effective subset of network edges for compressing the weight distribution across all edges in the network under the specified encoding scheme.
It is evident from its formulation that the global MDL method described above is simply a principled, nonparametric way to identify a global threshold on the weights for constructing the backbone. Therefore this method is naturally limited for constructing network backbones when node strengths are heterogeneous—this is a major motivation for local methods such as the Disparity or Polya Filter serrano2009extracting ; marcaccioli2019polya . In such cases, the backbone $G^{(b)}$ is likely to be highly fragmented and have many degree-zero nodes as a result of weakly interacting nodes losing all of their adjacent edges. However, when node strengths are highly homogeneous, the global MDL backboning method may still be statistically justified from the perspective of data compression (we will see this in Sec. III). Therefore, for any backboning task one should use both the global method of this section and the local method presented in the next section, then compare their minimum description length values to choose the optimal MDL backbone for the dataset $G$.
II.2 Local Network Backboning Algorithm
To construct a local network backboning objective, we can apply a similar encoding scheme as above but at the level of individual node neighborhoods, which corrects for heterogeneity in the overall strengths of the nodes in a similar manner as other local backboning methods serrano2009extracting ; marcaccioli2019polya . For simplicity of presentation, we will define each node neighborhood $\partial_{i}$ as the set edges with node $i$ as the source node. But the method can be applied in an identical manner by considering all edges in which $i$ is the target node (i.e. considering in-neighborhoods instead of out-neighborhoods). For undirected networks, this distinction will make no difference since the edge directions are duplicated, so we can just use the out-neighborhoods. Since the method described has a simple closed form description length, one can choose the backbone to be constructed based on in- or out-neighborhoods based on which provides better data compression according to the final description length value.
Let $\partial_{i}=\{(i,j):w_{ij}>0\}$ be the neighborhood of node $i$ (in terms of edges), with $k_{i}=|\partial_{i}|$ and $s_{i}=\sum_{(i,j,w_{ij})\in\partial_{i}}w_{ij}$ the degree and strength of node $i$ respectively. We also let $k_{i}^{(b)}$, $s_{i}^{(b)}$, and $\partial_{i}^{(b)}$ be defined analogously for the backbone network $G^{(b)}$, such that $\partial_{i}^{(b)}=\partial_{i}\cap G^{(b)}$ is the neighborhood of node $i$ in the backbone $G^{(b)}$. Transmitting the network $G$ is equivalent to transmitting $\partial_{i}$ for all nodes $i$, and we can formulate our objective in this way by using the same encoding as in the global case to transmit $\partial_{i}$ using $\partial_{i}^{(b)}$ for all nodes $i$ separately. This simply requires matching up the neighborhood level variables $\{s_{i},k_{i},s_{i}^{(b)},k_{i}^{(b)}\}$ with the global variables $\{W,E,W^{(b)},E^{(b)}\}$ in the description length of Eq. II.1.
Given this correspondence with the global method, the information required to transmit the neighborhood $\partial_{i}$ using the backbone neighborhood $\partial^{(b)}_{i}$ is given by
$\mathcal{L}^{(\text{neig})}(\partial_{i}^{(b)})$ $=\log(s_{i}-k_{i}+1)+\log(s_{i}+1)$
$+\log(k_{i}+1)+\log{k_{i}\choose k_{i}^{(b)}}$ (4)
$+\log{s_{i}^{(b)}-1\choose k_{i}^{(b)}-1}+\log{s_{i}-s_{i}^{(b)}-1\choose k_{i}-k_{i}^{(b)}-1}.$
The total description length for the local MDL backboning method is then given by
$\mathcal{L}^{(\text{local})}(G^{(b)})=\sum_{i=1}^{N}\mathcal{L}^{(\text{neig})}(\partial_{i}^{(b)}),$ (5)
where $N$ is the total number of nodes in the network.
Since the description length is analogous to the global description length of Eq. II.1, we can use a similar argument as in Sec. II.1 to show that in the regime $s_{i},k_{i}\gg 1$ the neighborhood-level description length $\mathcal{L}^{(\text{neig})}(\partial_{i}^{(b)})$ is monotonically decreasing in $s_{i}^{(b)}$ for any fixed value of $k_{i}^{(b)}$, so long as $\frac{s_{i}^{(b)}}{s_{i}}\geq\frac{k^{(b)}_{i}}{k_{i}}$. This implies that a greedy algorithm will also be approximately optimal for constructing the node neighborhoods $\partial_{i}^{(b)}$ within the backbone, and an efficient algorithm for constructing the entire backbone is as follows:
  1. 1.
    Sort the edges $(i,j,w_{ij})\in G$ in decreasing order of weight $w_{ij}$. Initialize the weight-ordered neighborhood $\partial_{i}=\{\}$ for each node $i$.
  2. 2.
    Iterate over the ordered list $G$ and for each edge $(i,j,w_{ij})$, add the edge to $\partial_{i}$. Then, for each node $i=1,...,N$, perform steps 3 - 5 below.
  3. 3.
    Initialize an empty backbone neighborhood $\partial^{(b)}=\{\}$ with $k_{i}^{(b)}=s_{i}^{(b)}=0$ and initialize an array DL to store the description length value for each value of $k_{i}^{(b)}$. Compute $\text{DL}[k_{i}^{(b)}=0]$ using Eq. II.2.
  4. 4.
    Iterate through the ordered edges in $\partial_{i}$, and for each edge $(i,j,w_{ij})$: increment $k_{i}^{(b)}\mathrel{+}=1$ and $s_{i}^{(b)}\mathrel{+}=w_{ij}$ and compute $\text{DL}[k_{i}^{(b)}]$ using Eq. II.2.
  5. 5.
    After adding the last edge from $\partial_{i}$, identify the value $k_{i,\text{local}}^{(b)}=\text{argmin}[\text{DL}]$ for which the description length was minimized, and set the MDL-optimal backbone neighborhood $\partial^{(b)}_{i,\text{local}}$ to be the first $k_{i,\text{local}}^{(b)}$ edges in the ordered $\partial_{i}$.
The above algorithm considers out-neighborhoods $\partial_{i}$, but an analogous procedure holds for in-neighborhoods if we swap $i$ and $j$ in the neighborhood, strength, and degree updates. As with the global method, the bottleneck of this local backboning algorithm is the initial step of ordering the edges by weight, so the method incurs a total runtime complexity of $O(E\log E)$. This is comparable to the complexity of the Disparity Filter serrano2009extracting in practice and more algorithmically efficient than other existing local backboning methods marcaccioli2019polya ; foti2011nonparametric .
We can see that through its information theoretic formulation, the local backboning method applies a weight threshold that adapts to the heterogeneity within each node neighborhood $\partial_{i}$. This will allow it to provide a more compressive backbone for networks with heterogeneous node strengths $s_{i}$ than the global method described in the previous section. To determine which method provides superior compression for a given network $G$, we can run each algorithm and compute the final description length values $\mathcal{L}^{(\text{local})}(G^{(b)}_{\text{local}})$ and $\mathcal{L}^{(\text{global})}(G^{(b)}_{\text{global}})$ using Eq. 5 and Eq. II.1 respectively to identify the lower of the two description lengths.
It is worth noting that one can guarantee that the local MDL backbone constructed using the objective in Eq. 5 is a single (weakly) connected component if all nodes have at least one edge in their neighborhood in the original graph (in the direction considered in the MDL objective). The cases $\{s_{i}^{(b)},k_{i}^{(b)}\}=\{0,0\}$ and $\{s_{i}^{(b)},k_{i}^{(b)}\}=\{s_{i},k_{i}\}$ give equal contributions to the description length, so if the optimal backbone sets $\partial^{(b)}_{i}=\{\}$ we can equivalently set $\partial^{(b)}_{i}=\partial_{i}$ and obtain the same description length. For the experiments in Sec. III we let $\partial^{(b)}_{i}=\{\}$ for these cases, but our code implementation allows for either option. We will see that regardless of this choice the local backbones tend to be more well connected than the global backbones for the networks studied.
It is also useful to construct a normalized inverse compression ratio $\eta\in[0,1]$ to determine the extent to which we can compress a network using its MDL-optimal backbones relative to a naive encoding where we do not use any backbone as an intermediate step (equivalent mathematically to setting $G^{(b)}=\{\}$ as the empty graph). The ratio we use here is given by
$\eta^{(\text{global/local})}=\frac{\mathcal{L}^{(\text{global/local})}(G_{\text{global/local}}^{(b)})}{\max[\mathcal{L}^{(\text{global})}(\{\}),\mathcal{L}^{(\text{local})}(\{\})]},$ (6)
where $G_{\text{global/local}}^{(b)}$ is the MDL-optimal backbone obtained by the global (local) method when computing $\eta^{(\text{global})}$ ($\eta^{(\text{local})}$). We can also choose to normalize Eq. 6 by the description length corresponding to the method of interest (global or local) rather than the maximum, but the max normalization is nice in practice because it allows us to compare (a) the absolute description lengths of the global and local methods and (b) their compression relative to the most naive encoding all in a single measure.
Fig. 1 shows a diagram of both the global and local backboning methods for a small synthetic example network. Code implementing these algorithms can be found at https://github.com/aleckirkley/mdl-network-backbones.
II.3 Bayesian Generalization
It is well known that MDL methods are often equivalent to maximum a posteriori (MAP) estimation in Bayesian generative models mackay2003information . This is also the case for the global and local MDL backboning procedures described in Sec. II.1 and Sec. II.2.
Consider a Bayesian generative model for the weights $\{w_{ij}\}$ of the network $G$ when its edge positions are known. We first assign each edge $(i,j)\in G$ to the (global) backbone $G^{(b)}$ with probability $\pi_{b}$ and to the set of non-backbone edges with probability $(1-\pi_{b})$. This is equivalent to specifying an indicator vector $\bm{b}$ for existence in the backbone such that $b_{ij}=1$ if $(i,j)\in G^{(b)}$ and 00 otherwise. This gives us a prior distribution over the backbone $\bm{b}$ of
$P(\bm{b}|\pi_{b})=\prod_{(i,j)\in G}\pi_{b}^{b_{ij}}(1-\pi_{b})^{1-b_{ij}}.$ (7)
Now assign a weight $w_{ij}$ to the edge $(i,j)$ that is drawn from a geometric distribution with a success parameter $\theta_{b_{ij}}$ that depends on whether we are in the backbone or not (i.e. if $b_{ij}=0$ or $b_{ij}=1$). This allows us to distinguish the weights in the backbone from those outside of it through a mixture model over the edge weights. (There are then two independent parameters, $\theta_{0}$ and $\theta_{1}$.) The likelihood of the weighted network $G$ and the backbone assignment $\bm{b}$ given the model parameters can then be written as
$P(G,\bm{b}|\theta_{1},\theta_{0},\pi_{b})$ $=P(G|\theta_{1},\theta_{0},\bm{b})P(\bm{b}|\pi_{b})$
$=\prod_{(i,j)\in G}[\pi_{b}(1-\theta_{1})^{w_{ij}-1}\theta_{1}]^{b_{ij}}$
×$\times[(1-\pi_{b})(1-\theta_{0})^{w_{ij}-1}\theta_{0}]^{1-b_{ij}}$
$=\pi_{b}^{E^{(b)}}(1-\pi_{b})^{E-E^{(b)}}$ (8)
×$\times\theta_{1}^{E^{(b)}}(1-\theta_{1})^{W^{(b)}-E^{(b)}}$
×$\times\theta_{0}^{E-E^{(b)}}(1-\theta_{0})^{(W-W^{(b)})-(E-E^{(b)})},$
where we have parameterized the geometric distribution so as to enforce $w_{ij}\geq 1$.
Putting a uniform prior $P(\theta_{1},\theta_{0},\pi_{b})=1$ on the model parameters, we can marginalize over these parameters on the unit interval $[0,1]$ to form the marginal joint likelihood
[…]
$=B(E^{(b)}+1,E-E^{(b)}+1)$ (9)
$\times$ $B(E^{(b)}+1,W^{(b)}-E^{(b)}+1)$
$\times$ $B(E-E^{(b)}+1,(W-W^{(b)})-(E-E^{(b)})+1),$
where
$B(x,y)=\frac{(x-1)!(y-1)!}{(x+y-1)!}$ (10)
is the Beta function.
Maximum a posteriori (MAP) estimation with this model then aims to solve the following optimization problem to identify the optimal backbone $\bm{\hat{b}}$
$\bm{\hat{b}}$ $=\operatorname*{arg\,max}_{\bm{b}}P(\bm{b}|G)$
$=\operatorname*{arg\,min}_{\bm{b}}\left\{-\log P^{(\text{global})}(G,\bm{b})\right\}.$ (11)
Substituting in Eq. II.3, some algebra shows that
$-\log P^{(\text{global})}(G,\bm{b})$ $=\mathcal{L}^{(\text{global})}(G^{(b)})$ (12)
$~{}~{}~{}~{}~{}~{}+O(\log W+\log E),$
where $\mathcal{L}^{(\text{global})}(G^{(b)})$ is given by Eq. II.1 and the second term contains terms of size less than or equal to $\log W+\log E$. Matching terms with Eq. II.1, we can see that the global description length exceeds $\log W+\log E$ by at least $\log{W-1\choose E-1}$, and so in the limit of $W,E\gg 1$ the fractional error between $-\log P^{(\text{global})}(G,\bm{b})$ and $\mathcal{L}^{(\text{global})}(G^{(b)})$ becomes vanishingly small. In other words, MAP estimation with this Bayesian model is asymptotically equivalent to the MDL minimization problem with the objective in Eq. II.1.
We can generalize this Bayesian model to the neighborhood level for generating the edge weights in $\partial_{i}$, allowing us to make a comparison with the local MDL objective in Eq. 5. In this model, each node $i$ has its own probability $\pi_{b}(i)$ that one of its adjacent edges $(i,j)\in\partial_{i}$ is in the backbone $G^{(b)}$. Similarly, each node $i$ has its own success rates $\theta_{1}(i)$ and $\theta_{0}(i)$ for the geometric distribution over edge weights. We can then compute the joint marginal likelihood $P^{(\text{local})}(G,\bm{b})$ as before by integrating out the parameters $\{\pi_{b}(i),\theta_{1}(i),\theta_{0}(i)\}$ for each node neighborhood $\partial_{i}$ and taking a product over the independent neighborhoods. The result is
$P^{(\text{local})}(G,\bm{b})$ $=\prod_{i=1}^{N}P(\partial_{i},\{b_{ij}\}_{(i,j)\in\partial_{i}})$
$=\prod_{i=1}^{N}B(k_{i}^{(b)}+1,k_{i}-k_{i}^{(b)}+1)$ (13)
$\times$ $B(k_{i}^{(b)}+1,s_{i}^{(b)}-k_{i}^{(b)}+1)$
$\times$ $B(k_{i}-k_{i}^{(b)}+1,(s_{i}-s_{i}^{(b)})-(k_{i}-k_{i}^{(b)})+1),$
where $P(\partial_{i},\{b_{ij}\}_{(i,j)\in\partial_{i}})$ is the local marginal joint likelihood of the neighborhood $\partial_{i}$ and its backbone edges $\{b_{ij}\}_{(i,j)\in\partial_{i}}=\partial^{(b)}_{i}$ after integrating over $\{\pi_{b}(i),\theta_{1}(i),\theta_{0}(i)\}$ with a uniform prior. (As before, the local model can be directly obtained from the global model by mapping the variables $\{W,E,W^{(b)},E^{(b)}\}\to\{s_{i},k_{i},s_{i}^{(b)},k_{i}^{(b)}\}$ and treating each neighborhood $\partial_{i}$ separately.) We can then a similar substitution as above to find
$-\log P^{(\text{local})}(G,\bm{b})$ $=\mathcal{L}^{(\text{local})}(G^{(b)})$ (14)
$+O\Big{(}\sum_{i=1}^{N}[\log s_{i}+\log k_{i}]\Big{)},$
with $\mathcal{L}^{(\text{local})}(G^{(b)})$ given by Eq. 5 and the Big-O notation used similarly as above.
Matching terms with Eq. II.2, we can see that the neighborhood-level description length exceeds $\log s_{i}+\log k_{i}$ by at least $\log{s_{i}-1\choose k_{i}-1}$, and so in the limit of $s_{i},k_{i}\gg 1$ the fractional error between $-\log P^{(\text{local})}(G,\bm{b})$ and $\mathcal{L}^{(\text{local})}(G^{(b)})$ becomes vanishingly small. Therefore MAP estimation with this Bayesian model is asymptotically equivalent to the MDL minimization problem with the objective in Eq. 5.
Generalizing beyond geometric weight distributions, we can formulate a flexible fully Bayesian generative model for global network backbones as follows, starting with an unweighted network:
  1. 1.
    Draw an inclusion probability $\pi_{b}$ according to a Beta prior $P(\pi_{b})$, which may be uniform. Assign each edge $(i,j)$ to the backbone $G^{(b)}$ with probability $\pi_{b}$ to construct the indicator vector $\bm{b}$.
  2. 2.
    Choose a discrete weight distribution $P(w_{ij}|\theta_{b_{ij}})$. Draw $\theta_{0},\theta_{1}$ from a factorizable prior $P(\theta_{0},\theta_{1})=P(\theta_{0})P(\theta_{1})$ that is conjugate to $P(w_{ij}|\theta_{b_{ij}})$ and may be uniform. Set the $\theta$ that gives a higher mean weight as $\theta_{1}$ if generating a high weight backbone.
  3. 3.
    Generate the weight $w_{ij}$ for each existing edge $(i,j)$ according to $P(w_{ij}|\theta_{b_{ij}})$.
We can also formulate a Bayesian generative model for local network backbones as follows:
  1. 1.
    For each neighborhood $\partial_{i}$, draw an inclusion probability $\pi_{b}(i)$ according to a Beta prior $P(\pi_{b}(i))$, which may be uniform. Assign each edge $(i,j)\in\partial_{i}$ to the backbone $G^{(b)}$ with probability $\pi_{b}(i)$ to construct the indicator vector $\bm{b}$.
  2. 2.
    Choose a discrete weight distribution $P(w_{ij}|\theta_{b_{ij}}(i))$ for each neighborhood $\partial_{i}$. Draw $\theta_{0}(i),\theta_{1}(i)$ from a factorizable prior $P(\theta_{0}(i),\theta_{1}(i))=P(\theta_{0}(i))P(\theta_{1}(i))$ that is conjugate to $P(w_{ij}|\theta_{b_{ij}}(i))$. The distributions $P(w_{ij}|\theta_{b_{ij}}(i))$ may be different for different nodes if desired.
  3. 3.
    Generate the weight $w_{ij}$ for each existing edge $(i,j)$ according to $P(w_{ij}|\theta_{b_{ij}}(i))$.
Both of these frameworks permit nonparametric MAP estimation by minimizing the description length $-\log P(G,\bm{b})$ over labellings (backbones) $\bm{b}$. The joint marginal likelihood $P(G,\bm{b})$ can be computed by integrating out the model parameters which have the appropriate conjugate priors for easy marginalization. This general framework also allows for model selection among both weight distributions and global or local schemes to determine the optimal backbone for a given network. However, depending on the weight distribution one chooses, the optimization may not be as simple as the greedy scheme described here. In these cases one can minimize over $\bm{b}$ using discrete MCMC methods or perhaps with continuous optimization methods for a relaxation of $\bm{b}$ to a real-valued vector.
III Results
III.1 Synthetic Backbone Reconstruction
Figure 2: Reconstruction of planted backbone structure in synthetic network data. (a) Network Jaccard similarity (Eq. 15) between the planted and inferred backbones $G^{(b)}_{planted}$ and $G^{(b)}_{MDL}$, using the global and local MDL objectives in Eq.s II.1 and 5 respectively. The difficulty of the reconstruction task is varied by changing the noise parameter $\gamma=\theta_{1}/\theta_{0}$, which is the ratio of the means for the non-backbone and backbone weight distributions, and the degree $k$ of the nodes in the underlying random regular directed graph with $N=100$ nodes. The synthetic networks are drawn from the global Bayesian generative model with geometric weights discussed in Sec. II.3. (b) Reconstruction performance for synthetic networks drawn from the local Bayesian generative model with geometric weights discussed in Sec. II.3. (c) Inverse compression ratio (Eq. 6) for the same set of experiments as in panel (a). (d) Inverse compression ratio for the same experiments as in panel (b). Markers indicate averages over 100 simulations and error bars represent one standard error in the mean.
As a first set of experiments to test our backboning methods, we examine the ability of these methods to reconstruct planted backbone structure in synthetic data simulated from the Bayesian generative models described in Sec. II.3. In the experiments we first generate an unweighted, directed graph with a fixed number of out-edges $k$ for each node $i$. The focus of the MDL backboning methods is on the edge weight distributions, and degree fluctuations simply provide noise in the reconstruction results so degrees are set to be uniform across nodes. We then generate the weights for this network using the global and local generative schemes described in Sec. II.3, with the priors $P(\pi_{b})$ and $P(\theta_{0})$ (for the global generative model) or $P(\theta_{0}(i))$ (for the local generative model) uniform over the range $[0,1]$ and $P(w_{ij}|\theta_{b_{ij}})$ or $P(w_{ij}|\theta_{b_{ij}}(i))$ set to a geometric distribution. As a measure of how “noisy” the backbone is—in other words, how similar the backbone weights and non-backbone weights are—we fix a parameter $\gamma=\theta_{1}/\theta_{0}$ for each simulation which is equivalent to the ratio of the mean non-backbone edge weight and mean backbone edge weight. We therefore set $\theta_{1}=\gamma\theta_{0}$ for each simulation to fix the ratio of the mean weights.
We test the reconstruction and compression capability for the global and local MDL backboning algorithms in Sec.s II.1 and II.2 using two measures. The first is the Jaccard similarity index between the planted backbone $G^{(b)}_{\text{planted}}$ generated from the model and the inferred backbone $G^{(b)}_{\text{MDL}}$ for each method. Since the networks $G$ are represented as edge sets, the Jaccard similarity index between a network $G_{1}$ and a network $G_{2}$ can be computed as
[…]
The Jaccard similarity in this case tells us how well the global and local MDL backboning algorithms can recover planted structure from their own (approximate) generative models, and falls in the range $[0,1]$ with 00 indicating an inferred empty or complete graph and $1$ indicating perfect recovery. There are many alternative suitable measures for computing the graph similarity felippe2024network , but in practice they give the same qualitative trends as the Jaccard similarity for this example. The second measure we use to evaluate the methods is the inverse compression ratio $\eta$ (Eq. 6), which tells us how well the methods compress the synthetic network data.
Fig. 2 shows the results of our reconstruction experiments, which were run for random directed regular graphs with $N=100$ nodes, degree $k\in\{5,20,100\}$, and varying noise levels $\gamma\in[0,1]$. Both the global (left column of panels) and local (right column of panels) generative models were used, to determine the extent to which each backboning method is robust under model misspecification. We can see from Fig. 2(a) that, as expected, the reconstruction performance becomes worse as we increase the noise level $\gamma$. At $\gamma\approx 10^{-3}$ we observe near perfect backbone recovery for both methods in the high degree regime, but at $\gamma\approx 1$ we see that neither inferred backbone has any significant shared structure with the planted backbone. We can see that the global method is insensitive to degree $k$ while the local method is sensitive to $k$, since lower $k$ will result in greater weight fluctuations within the node neighborhoods but less significant fluctuations at the global level since $E=500$ even for the lowest degree value ($k=5$) studied. We also see that, as expected, the global MDL method has better performance in the low noise regime than the local MDL method, since it is (approximately) the Bayes-optimal algorithm. However, for higher levels of noise $\gamma$ we see that the local backboning method achieves better performance, which improves for smaller degrees $k$. This is because in this regime the global backboning method tends to infer very sparse backbones, but the local method can still find backbones with a moderate level of connectivity due to considering within-neighborhood weight fluctuations rather than global weight fluctuations.
In Fig. 2(b) we show the reconstruction performance for the same set of experiments but for the local generative model, which allows the model parameters $\pi_{b}(i),\theta_{0}(i),\theta_{1}(i)$ to vary for each node neighborhood $\partial_{i}$. This causes fluctuations in the weight distributions across nodes, making it more challenging in general to infer the correct backbone structure. Indeed, we can observe in Fig. 2(b) that both methods exhibit a modest drop in reconstruction performance, with the local MDL method now outperforming the global MDL method in the low noise regime for high degrees. Neither method is capable of perfect reconstruction for low noise levels in this more challenging task, but both can still recover the planted backbone with reasonable accuracy.
We plot the inverse compression ratio (Eq. 6) versus the noise level $\gamma$ in Fig. 2(c) and Fig. 2(d) for the global and local generative models respectively. We see that for the global generative model the global MDL method outperforms the local MDL method, and for the local generative model the local MDL method outperforms the global MDL method. This is consistent with the global and local methods being approximately Bayes-optimal for the global and local Bayesian models respectively, as discussed in Sec. II.3. In both cases substantial compression is possible using both methods, and the networks are more compressible for smaller degrees $k$.
The results in Fig. 2 are consistent with the arguments in Sec. II.3 for the correspondence between the proposed MDL methods and Bayesian geometric generative models of network weights. They also demonstrate the capability for these methods to identify meaningful backbone structure and provide considerable compression of network data in the presence of noise. In the next section we compare the MDL methods against existing methods on synthetic data with various levels of planted global and local homogeneity in the edge weights.
III.2 Comparison on Synthetic Networks
Figure 3: Concentration of weight along different backbones. Top row: Fraction of edges retained in the backbone, $E^{(b)}/E$, versus the (a) average node strength $W/N$, (b) level of homogeneity $h_{\text{str}}$ in the node strength distribution, and (c) level of homogeneity $h_{\text{neig}}$ in the weights within each node neighborhood. Bottom row: Panels (d)-(f) plot the fraction of weight retained in the backbone, $W^{(b)}/W$, against the same parameters. Markers indicate averages over 10 simulations from the synthetic network model described in Sec. III.2, and error bars represent one standard error in the mean.
Figure 4: Global connectivity of different backbones. Top row: Fraction of non-isolated nodes retained in the backbone versus the (a) average node strength $W/N$, (b) level of homogeneity $h_{\text{str}}$ in the node strength distribution, and (c) level of homogeneity $h_{\text{neig}}$ in the weights within each node neighborhood. Bottom row: Panels (d)-(f) plot the relative reachability (Eq. 16) against the same parameters. Markers indicate averages over 10 simulations from the synthetic network model described in Sec. III.2, and error bars represent one standard error in the mean.
Figure 5: Compression and node strength discrepancies among different backbones. Top row: Inverse compression ratio (Eq. 6) versus the (a) average node strength $W/N$, (b) level of homogeneity $h_{\text{str}}$ in the node strength distribution, and (c) level of homogeneity $h_{\text{neig}}$ in the weights within each node neighborhood. Bottom row: Panels (d)-(f) plot the Hellinger distance (Eq. 17) between the original strength distribution and the backbone strength distribution against the same parameters. Markers indicate averages over 10 simulations from the synthetic network model described in Sec. III.2, and error bars represent one standard error in the mean.
Figure 6: Global and local backbones across a real-world network corpus. Global and local MDL backbones were inferred using the objectives of Eq. II.1 and Eq. 5 respectively, for the set of real networks described in Sec. III.3. We plot (a) the fraction of initial edges retained in the local MDL backbone and global MDL backbone and (b) the inverse compression ratio (Eq. 6) for each backbone examined. We only include examples for which both the global and local MDL methods inferred a non-empty backbone, noting the empty backbones in the axis labels. In both panels we plot the line $y=x$ for reference and color the examples for which the global (local) backbone returned a higher value with blue (red). We also indicate with different markers the domain of each example network.
As discussed in the Introduction, the MDL methods we propose in this paper are fully nonparametric, in contrast with existing methods for network backboning which either constrain the topology of the backbone to a specific structure such as a tree, require the specification of the desired number of edges $E^{(b)}$ for the backbone, or require a significance level at which to retain edges under some null model for the weights serrano2009extracting ; foti2011nonparametric ; marcaccioli2019polya ; yassin2023evaluation . This makes direct comparison with existing methods challenging.
We therefore select a few popular existing methods for global and local backboning for comparison, with parameters fixed as in the original studies. The first method we use for comparison is the Disparity Filter serrano2009extracting , which is a principled inference-based method for local backboning that selects edges below a pre-specified significance level $\alpha$ within each node neighborhood for the backbone. The null model used for comparison is one in which the weight assignment is uniform conditioned on the node degree and strength. For more direct comparison with our own local method, for one baseline we fix $E^{(b)}=|G^{(b)}_{\text{local}}|$ as the number of edges to retain in the Disparity Filter and set $\alpha$ to obtain the desired number of edges. For the other Disparity Filter baseline, we set the p-value threshold to the common value of $\alpha=0.05$. We constrain the method to focus on out-neighborhoods for these synthetic examples.
For the other baseline, we compare against a popular global backboning method called the High Salience Skeleton grady2012robust . This method computes the saliency of a link based on its occurrence frequency in the shortest path trees rooted at each node, giving a succinct global view of edge importance. It is observed that the saliency distribution of links in real networks is highly bimodal, so the skeleton is not sensitive to the specific choice of the saliency cutoff for the backbone, which is chosen to be the center of the saliency distribution. The same methodology is applied here for the baseline. Due to its computation of the shortest path tree for each node, the High Salience Skeleton can become too computationally expensive for large networks, so for $N>10000$ we randomly sample the shortest path trees for $10000$ nodes to estimate the link saliency.
The synthetic networks used for the comparison experiments were generated using the following procedure. First, as in Sec. III.1, an unweighted network of $N$ nodes is generated as a $k$-regular random directed graph with the specified out-degree $k$. Given an input parameter $W$ for the total edge weight, all edges are then assigned weight $1$ and the excess weight $W-Nk$ is distributed across the nodes’ out-neighborhoods using a symmetric Dirichlet-Multinomial distribution with concentration parameter $h_{\text{str}}$. This allows us to tune the level of homogeneity in the out-strengths $\{s_{i}\}$ across nodes with a single parameter $h_{\text{str}}$: $h_{\text{str}}\to 0$ places all the excess edge weight in a single out-neighborhood $\partial_{i}$ to give a highly heterogeneous strength distribution, and $h_{\text{str}}\to\infty$ distributes all the weight equally across nodes to give a perfectly homogeneous strength distribution. We then distribute the total excess edge weight $s_{i}-k$ among the out-edges in the neighborhood $\partial_{i}$ using another symmetric Dirichlet-Multinomial distribution, this time with concentration parameter $h_{\text{neig}}$. $h_{\text{neig}}\to 0$ places all the excess edge weight on a single edge $w_{ij}$ to give a highly heterogeneous weight distribution in each neighborhood, and $h_{\text{neig}}\to\infty$ distributes all the weight equally across edges to give a perfectly homogeneous strength distribution. The synthetic networks therefore depend on the four parameters $\{N,W,h_{\text{str}},h_{\text{neig}}\}$, which we vary in our experiments.
We compare backboning methods using multiple indicators, similar to the methodology used in yassin2023evaluation . We compare the fraction of edges $E^{(b)}/E$ and total weight $W^{(b)}/W$ appearing in the backbone to get a sense of how sparse the inferred backbones are with each method. We also compute the inverse compression ratio (Eq. 6) using the global description length (Eq. II.1) for our global MDL method and the High Salience Skeleton, and the local description length (Eq. 5) for our local MDL method and the Disparity Filter variants to see how well each method compresses the data according to the global and local encodings. To examine the extent to which the backbones maintain global connectivity we compute the fraction of nodes $N^{(b)}/N$ with non-zero degree in the backbone and the relative reachability $R$, defined by
$R(G^{(b)})=\frac{\text{\# pairs $i,j$ with directed path $i\to j$ in $G^{(b)}$}}{\text{\# pairs $i,j$ with directed path $i\to j$ in $G$}}.$ (16)
Since $R(G^{(b)})$ can become computationally prohibitive for large networks, for $N>10000$ we randomly sample a subgraph of $10000$ nodes to estimate the reachability. We also determine the extent to which the strength distribution has been preserved across nodes by computing the Hellinger distance between the strengths $\{s_{i}\}_{i=1}^{N}$ in the original network $G$ and the strengths $\{s^{(b)}_{i}\}_{i=1}^{N}$ in each backbone $G^{(b)}$ using
$D_{\text{str}}(G,G^{(b)})=\sqrt{\frac{1}{2}\sum_{i=1}^{N}(\sqrt{p_{i}}-\sqrt{q_{i}})^{2}},$ (17)
where $p_{i}=s_{i}/W$ and $q_{i}=s_{i}^{(b)}/W^{(b)}$.
In Fig. 3 we plot the fraction of edges (top row) and weight (bottom row) retained in the different backbones of the different backboning methods versus the synthetic model parameters. Unless otherwise specified as an independent variable, in all simulations we set $N=1000$, $W/N=1000$, $k=50$, and $h_{\text{str}}=h_{\text{neig}}=0.1$. There are a few major trends we can observe here. The first is that when $h_{\text{neig}}=0.1$ is held constant at a moderate value (first two columns in Fig. 3), both MDL methods consistently retain a similar number of edges, which is the highest among all methods (an average of around $20\%$ of all edges). We can also see that the High Salience Skeleton consistently retains the fewest edges (an average of around $2\%$ of all edges) in these cases, while the Disparity Filter with an $\alpha=0.05$ significance level finds an intermediate number of backbone edges (an average of around $8\%$ of all edges). We can also see that the local MDL method consistently retains a higher total backbone weight $W^{(b)}$ for the same number of edges $E^{(b)}$ as the Disparity Filter when the latter is set to have the same edge count. The three local backboning methods are highly sensitive to variations in the heterogeneity of weights within the node neighborhoods $h_{\text{neig}}$, while the global MDL method and the High Salience Skeleton are less sensitive to these fluctuations. The local backboning methods all have a sudden drop in the number of edges and total weight they retain at $h_{\text{neig}}\approx 1$, which is the point at which the within-neighborhood variance in the edge weights drops significantly according to this Dirichlet concentration parameter. (This low level of weight homogeneity within neighborhoods leaves little statistical evidence to retain any edges in the backbone.) The local MDL method is the least sensitive to these fluctuations among the local backboning methods.
While the size of the backbones is an important factor for efficient downstream graph computations, it is also important that the sparsified networks retain their global connectivity. In the top row of Fig. 4 we plot the fraction of non-isolated nodes in the backbone (i.e. nodes with degree 00) versus the three model parameters. We can observe that all backbones except the High Salience Skeleton have zero isolated nodes for a large portion of the parameter space in panels (a) and (b), with higher levels of node isolation for very low levels of strength homogeneity $h_{\text{str}}\lesssim 10^{-2}$ due to sparser backbones across the board. In panel (c) we again see strong fluctuations for the local backboning methods—particularly for the Disparity Filter at $\alpha=0.05$—with the MDL method again performing the best out of the local methods and the global MDL method outperforming the High Salience skeleton in terms of node connectivity. The bottom row shows similar connectivity patterns across methods but this time with respect to reachability (Eq. 16). Here the Disparity Filter has a slight edge over the local MDL method in most experiments for the same number of edges and consistently worse reachability for $\alpha=0.05$. The High Salience Skeleton has exactly zero reachabilty in all experiments, suggesting that although this method highlights important edges for global routing it may tend to construct backbones that are too sparse to maintain global connectivity in the network. All methods except the global MDL method perform poorly in the regime of high neighborhood weight homogeneity, becoming too sparse to have a high reachability value.
In Fig. 5 we plot a few additional metrics against the synthetic model parameters. In the top row, we can see that (as expected) the global MDL method is the most compressive among the global methods and the local MDL method is the most compressive among the local methods. Although the local methods tend to be more compressive in general, for high strength homogeneity (panel (b)) we find a slight compressibility advantage for the global method, suggesting that for networks with simple weight distributions one may favor the global MDL method over the local MDL method due to the additional model complexity of the latter. All methods exhibit worse compression as the weights become more homogeneous, since there is little statistical evidence for a separate backbone of high weight edges. In the bottom row we plot the Hellinger distance between the node strength distributions in the full network and the backbone (Eq. 17), which shows similar performance for the global MDL method, local MDL method, and Disparity Filter with $E_{Local}$, while the Disparity Filter at $\alpha=0.05$ has slightly worse performance (e.g. a higher discrepancy in the strength distributions). In panel (f), however, we see that the MDL methods are much more robust to neighborhood weight heterogeneity fluctuations than the other methods, maintaining relatively similar strength distributions in the inferred backbones as the original networks.
In Appendix A we plot the sizes of the inferred backbones versus the number of nodes in the original network (Fig. 8), finding a very weak relationship with this model parameter. We also show the runtime of the different backboning methods versus the network size, in Fig. 9. The slightly superlinear scaling of the MDL runtimes is consistent with the theoretical scaling of $O(E\log E)$ discussed in Sec. II.1 and Sec. II.2. We also observe a substantially higher computational complexity for the High Salience Skeleton due to the computation of the shortest path trees.
III.3 Comparison on Real Networks
Figure 7: Comparison of backbone metrics on real networks. The five backboning methods are compared using the same metrics as in Sec. III.2 on the real networks in the corpus described in Sec. III.3. In each of the first five panels we plot a matrix $\bm{B}(m)$ for a metric $m$ of interest such that $B_{ij}(m)$ is the fraction of real networks in the corpus for which $m_{i}\geq m_{j}$. High values $B_{ij}(m)$ indicate that the metric $m$ was higher for backboning method $i$ than backboning method $j$ in a large portion of real network instances. We plot the matrix $\bm{B}(m)$ for (a) the fraction of edges retained in the backbone, $m=E^{(b)}/E$; (b) the fraction of nodes with degree zero in the backbone; (c) the Hellinger distance (Eq. 17) between the strength distributions of the backbone and original network; (d) the fraction of weight retained in the backbone, $m=W^{(b)}/W$; and (e) the relative reachability (Eq. 16) between the backbone and original graph. In panel (f) we plot the average network Jaccard similarity (Eq. 15) between the backbones generated by each pair of methods.
We now compare the five backboning algorithms from Sec. III.2 on a large corpus of real network datasets coming from different application domains. We initially collected 69 networks from the Netzschleuder repository netz by retrieving all weighted networks with edge weights $w_{ij}\geq 1$ and less than $10^{7}$ edges. Non-integer-valued weights $w_{ij}$ were rounded to the nearest integer for analyses. As the methods in Sec. III.2 are all adaptable to both directed and undirected networks, we analyzed the 32 directed networks in the corpus using both edge directions separately (in-edges and out-edges) and analyzed the remaining 37 undirected networks with both edge directions simultaneously as described in Sec. II.1. Three directed networks and three undirected networks returned empty backbones for all five methods so were discarded from the results, leaving 58 directed and 34 undirected backbones constructed by each of the five methods. The networks in the final corpus represent a diverse array of domains, and (using the classification provided by netz ) in total we inferred 54 social network backbones, 11 informational network backbones, 9 biological network backbones, 7 technological network backbones, 6 transportation network backbones, and 5 economic network backbones using each of the five methods. The networks ranged in size from $E=91$ to $E=5,743,258$ with total weight ranging from $W=282$ to $W=2,541,576,441$ and average degree ranging from $\langle k\rangle=1.1$ to $\langle k\rangle=238.3$.
Our first experiment with this real-world network corpus examines the extent to which the backbones differ between the global and local MDL backboning procedures described in Sec.s II.1 and II.2. In Fig. 6(a) we plot the fraction of the original edges retained in the local MDL backbone versus the fraction retained in the global MDL backbone for each of the backbones examined. We only include examples for which both methods found a non-empty backbone, noting the number of violating examples in the axis labels. We can see that the global MDL method tends to find empty backbones in a large fraction of the examples studied (25/92), indicating that this method is frequently not able to compress real network structure any better than a naive edge transmission. On the other hand, only two networks failed to be compressed by the more flexible local MDL method. When both methods give nontrivial backbones, we can see that they return a similar number of edges, with the local MDL method returning more edges in many cases but neither method exceeding $E^{(b)}/E\approx 0.35$. In Fig. 6(b) we plot the inverse compression ratios (Eq. 6) for each method, which indicates that when the global MDL method does compress it does so quite well, even compressing roughly $\approx 90\%$ of the information in the network relative to a naive edge transmission in some cases. We see that the global method compresses better than the local method in many cases, demonstrating that a global threshold can in fact provide an effective backbone for many real networks. Neither panel indicates any clear differentiation among the backbone structure of networks based on domain.
In the next experiment we compare all five methods with respect to the metrics evaluated in Sec. III.2—specifically, the fraction of edges retained in the backbone; the fraction of nodes with degree zero in the backbone; the Hellinger distance (Eq. 17) between the strength distributions of the backbone and original network; the fraction of weight retained in the backbone, $m=W^{(b)}/W$; and the relative reachability (Eq. 16) between the backbone and original graph. We omit analysis of the inverse compression ratio since trivially the global and local MDL methods will always obtain better compression than their global and local counterparts, and the direct comparison of the two MDL methods is done in Fig. 6. Fig. 7(a)-(e) shows the results of these experiments through pairwise comparisons among all pairs of backboning methods for each of the five measures. We can see that in the real networks the methods are roughly ordered like MDL $\text{MDL Local}\approx\text{Disparity}(E_{Local})>\text{MDL Global}>\text{Disparity}(\alpha=0.05)>\text{High Salience}$ when considering the number of edges, the total weight, the fraction of isolated nodes, and the relative reachability of the inferred backbones. The reverse pattern is observed for the Hellinger distance, with MDL Local and $\text{Disparity}(E_{Local})$ performing better than the other methods on a substantial fraction of networks. MDL Local has a slight edge over $\text{Disparity}(E_{Local})$ with respect to the fraction of isolated nodes and the reachability, while $\text{Disparity}(E_{Local})$ has a slight edge for the Hellinger distance.
In Fig. 7(f) we plot the Jaccard similarity (Eq. 15) between the backbones generated by each pair of methods, averaged over all networks in the corpus. We observe very low levels of similarity among all pairs of backbones except for those generated by the MDL Local and $\text{Disparity}(E_{Local})$ methods, which have a moderately high average overlap of roughly $0.5$.
Overall, the results in Sec. III.2 and Sec. III.3 indicate that the local MDL method and the Disparity Filter set to the same number of backbone edges as the local MDL method tend to best preserve the original network structure in many cases. Meanwhile, the global MDL method often finds empty backbones and does not compress relative to a naive transmission scheme, but in the instances that it does compress it performs well across the metrics while retaining fewer edges and providing even better compression than the local method. Setting the Disparity filter to the typical significance level of $\alpha=0.05$ tends to be too conservative in some real networks, suggesting that this parameter needs to be carefully tuned in practice for different applications. Finally, the High Salience Skeleton is quite conservative in its backbone estimation, as it highlights only a few links that are the most significant for global network routing.
IV Conclusion
In this paper we develop a completely nonparametric framework for inferring the backbone of a weighted network which utilizes the minimum description length (MDL) principle to promote sparsity. Our method is adapted to infer both global and local network backbones, and is generalizable to any mixture of weight distributions over the backbone and non-backbone edges using a corresponding Bayesian generative model. We develop fast optimization schemes for our global and local MDL backboning objectives that are log-linear in the number of edges, allowing these principled methods to easily scale to networks with millions of edges. We compare our method with existing methods in a range of tasks on synthetic and real network data, showing that the proposed MDL methods are capable of substantially sparsifying a wide variety of networks while retaining meaningful global and local structural characteristics.
There are a number of potential avenues for future work extending our methods. As mentioned in Sec. II.3, our method is generalizable for inferring global and local network backbones under any mixture of weight distributions in the backbone and non-backbone. Here we study a model equivalent to a mixture of geometric distributions, but in future work it is important to examine other distributions which may provide better compression of networks with different levels of weight heterogeneity. One could also fully utilize the posterior distribution over backbones in this Bayesian framework to sparsify networks in a stochastic manner, similarly to some existing sparsification techniques spielman2008graph . Both of these extensions may incur a substantial additional computational burden for optimization and sampling compared to the fast greedy optimization performed here. One could also extend our framework to sparsify hypergraphs or other higher order networks with weight metadata. Finally, it may be possible to extend our method to infer sparse functional network backbones by exploiting regularities in dynamical information such as node or edge states.
Acknowledgments
The author acknowledges support from the HKU-100 Start Up Fund.
Appendix A Additional Comparisons on Synthetic Networks
In this Appendix we provide additional tests to compare the different backboning methods with the synthetic network model of Sec. III.2.
Figure 8: Sizes of backbones for different network sizes. (a) Fraction of edges retained in the backbone, (b) fraction of total weight retained in the backbone, and (c) fraction of non-isolated nodes in the backbone versus the number of nodes in the synthetic networks generated with the model in Sec. III.2. Markers indicate averages over 10 simulations from the synthetic network model described in Sec. III.2, and error bars represent one standard error in the mean.
Figure 9: Runtime scaling of different backboning methods. Runtime per backbone evaluation (in seconds) versus the number of nodes $N$ for the synthetic network experiments in Sec. III.2. The global and local MDL backbones are computed together, as well as the Disparity Filter backbones. Best-fit slope values $\beta$ for linear fits of the form $\log(\text{Runtime})=\beta\log(N)+C$ are shown alongside each method in the legend. Markers indicate averages over 10 simulations from the synthetic network model described in Sec. III.2, and error bars represent one standard error in the mean.